OpenStudy (firejay5):

Define a variable, write an equation, and solve the problem: A school conference room can seat a maximum of 83 people. The principal and two counselors need to meet with the school's student athletes to discuss eligibility requirements. If each student must bring a parent with them, what is the maximum number of students that can attend each meeting?

OpenStudy (firejay5):

I need help people

hartnn (hartnn):

ok,let there be 'x' number of students that can attend the meeting. so including their parents,there will be 2x people now including the princi and 2 counselors, we have total of 2x+1+2 people ,this cannot exceed 83 so to get x,we write 2x+1+2=83 so,2x=80 so x=40 number of students that can attend the meeting=40

hartnn (hartnn):

u got it @Firejay5 or still have doubts?

OpenStudy (anonymous):

let there b x students and their parents is also x x+x=2x principal and 2 councellers=3 maximum number of seat=83 therefore equation=2x+3=83 2x=80 x=40 maximum no of students attending=40